OpenStudy (anonymous):

I need help on this question If 3a + 2b = 7 and 2a + 2b = 9 , what is the value of ?

OpenStudy (anonymous):

3a + 2b = 7---(1) 2a + 2b = 9---(2) From (1), a = (7-2b)/3---(3) Sub. (3) into (2), 2[(7-2b)/3]+2b = 9 the solve it

OpenStudy (anonymous):

I have three more I'm having trouble with. If I ask for help one at a time, could you help me?

OpenStudy (anonymous):

of course :)